Outline

Contents

1 Objectives for 1.105.2 1

2 What is the kernel? 2

3 Compiling a Kernel 4 3.1 Getting the sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Configuring the kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Compiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 Installing the kernel 7 4.1 Installing the kernel itself . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.2 Make the initial RAM disk file . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3 Having Grub start the kernel . . . . . . . . . . . . . . . . . . . . . . . . . 8

1 Objectives for 1.105.2

Description of Objective

Candidates should be able to customize, build, and install a kernel and kernel loadable modules from source. This objective includes customizing the current kernel configuration, building a new kernel, and building kernel modules as appropriate. It also includes installing the new kernel as well as any modules, and ensuring that the boot manager can locate the new kernel and associated files (generally located under /boot, see objective 1.102.2 for more details about boot manager configuration).

Key files, terms, and utilities include:

/usr/src/linux/*
/usr/src/linux/.config
/lib/modules/kernel-version/*
/boot/*
make
make targets: config, menuconfig,
xconfig, oldconfig, modules, install,

modules_install dep
depmod

2 What is the kernel?

What is the kernel?

• The kernel consists of:

– the kernel itself:

∗ such as /boot/vmlinuz-2.6.12-1.1447_FC4smp

– The kernel modules:

∗ In /lib/modules/$(uname -r)

Kernel naming conventions

• In a name such as vmlinuz-2.6.12-1.1447_FC4smp, there are the following parts of the name that identify the kernel: major number: here 2

– In the Makefile, called VERSION minor number: here 6

– In the Makefile, called PATCHLEVEL revision: here 12

– In the Makefile, called SUBLEVEL vendor string: here -1.1447_FC4smp

– In the Makefile, called EXTRAVERSION

– Always change this in the top level makefile if you already have an existing

kernel with the same name whose modules you do not want to overwrite

• The value that you choose for these variables in the top level Makefile determines what you see when you run the program uname -r

• Consequently also determines the name of the modules directory.

Types of Kernel Files

• The main kernel file can be produced by make zImage or by make bzImage

– a zImage kernel is limited to about 508 kB in size and is loaded into lower memory

– zImage kernels are deprecated after 2.4.0-test3-pre3

    ∗ See Documentation/i386/boot.txt

– a bzImage kernel can be up to about 2.5 MB in size

– Both are compressed using gzip compression

– the ‘b’ in “bzImage” means “big” rather than indicating bzip2 compression – bzImage kernels are loaded into higher memory

Other Kernel Files in /boot

• System.map contains the addresses of kernel symbols

– Used by tools to interpret kernel error messages or OOPSes, to translate kernel addresses into names that mean more to us humans

– See http://www.dirac.org/linux/system.map/

• The initrd file, which is a compressed filesystem that is mounted as a ram disk

– It contains the drivers (kernel modules) that the kernel needs to access the hard disk.

– The memory used by the initial ram disk is freed up after the modules have been

loaded into the kernel

• It is nice to have the kernel .config configuration file handy so that the administrator knows how the kernel was built

Kernel Modules

• Kernel modules usually provide one of the following:

device driver: supporting a specific kind of hardware file system driver: supporting the ability to read/write different file systems system call extensions: most system calls are supported by the base kernel, but mod-

ules can extend or add system calls network driver: implement particular network protocols executable loader: support loading and executing additional executable file formats

3. Compiling a Kernel  4

Documentation

• The directory Documentation under the top level contains lots of documentation relating to many aspects of the kernel.

• The file Documentation/Configure.help provides help with configuration for pre-2.6 kernels.

3 Compiling a Kernel

3.1 Getting the sources

Getting the sources

$ lftp ftp://ftp.au.kernel.org/pub/linux/kernel/v2.6/
cd ok, cwd=/pub/linux/kernel/v2.6
lftp ftp.au.kernel.org:/pub/linux/kernel/v2.6> ls
-rw-rw-r--     1 ftp       ftp          12777 Dec 18  2003 ChangeLog-2.6.0
-rw-rw-r--     1 ftp       ftp         193569 Jan 09  2004 ChangeLog-2.6.1
-rw-rw-r--     1 ftp       ftp        1552868 Dec 25  2004 ChangeLog-2.6.10
-rw-rw-r--     1 ftp       ftp        1495678 Mar 03  2005 ChangeLog-2.6.11
-rw-r--r--     1 ftp       ftp           1221 Mar 09  2005 ChangeLog-2.6.11.1
...
-rw-rw-r--     1 ftp       ftp        4191691 Oct 19  2004 patch-2.6.9.gz
-rw-rw-r--     1 ftp       ftp             248 Oct 19  2004 patch-2.6.9.gz.sign
-rw-rw-r--     1 ftp       ftp             248 Oct 19  2004 patch-2.6.9.sign
drwxrwsr-x     2 ftp       ftp           8192 Dec 19  2003 pre-releases
drwxrwsr-x     4 ftp       ftp          28672 Sep 13 03:05 snapshots
drwxrwsr-x     4 ftp       ftp          24576 Sep 13 13:53 testing
lftp ftp.au.kernel.org:/pub/linux/kernel/v2.6>  mget linux-2.6.13.1.tar.bz2*
 38375702 bytes transferred in 746 seconds (50.2K/s)
Total 2 files transferred
lftp ftp.au.kernel.org:/pub/linux/kernel/v2.6> bye
$ tar xvjf linux-2.6.13.1.tar.bz2
drwxr-xr-x git/git             0 2005-09-10 12:42:58 linux-2.6.13.1/
-rw-r--r-- git/git        18691 2005-09-10 12:42:58 linux-2.6.13.1/COPYING
-rw-r--r-- git/git        89317 2005-09-10 12:42:58 linux-2.6.13.1/CREDITS
drwxr-xr-x git/git             0 2005-09-10 12:42:58 linux-2.6.13.1/Documentation/
...

Where to untar the source?

• Many people untar the source below /usr/src

• … but I prefer to untar it in a subdirectory below my home directory

• It is better to compile the code as a normal user rather than as root

– It is a good principal to do anything with the least privelege required

• I will call the first directory appears when we untar the code as the top level directory

– For example, if I did the untarring above in the directory ∼/src, then the top

         level directory is ∼/src/linux-2.6.13.1

3.2 Configuring the kernel

editing .config

• We next need to edit/create a file .config in the top level directory

• Could edit by hand, but easy to make a mistake

• We call make with one of the four targets:

config this is a method I have not used for years. It does not allow you to go back: you can only move forward, answering questions

menuconfig this gives you a nice text curses-based screen that allows you to navigate through the choices as you wish

xconfig on 2.4 kernels, gives a nice Tk interface, and on 2.6 kernels gives a program called qconf, which on my system is linked to a qt library.

oldconfig this allows you to easily update an existing .config file, answering the configuration questions only for new options which are in the new source code, but which were not covered in the old .config file.

Answering the questions

• For each configuration option, we may be presented with the options

y yes: means compile this right into the base kernel

m module: means compile this as an external module that can be loaded into the kernel when it is needed

– It doesn’t hurt to compile lots of modules, even though you don’t need them, except that:

∗ it takes more time to compile,

∗ the chance of finding an error in the source code is increased, and

∗ the modules directory will take more hard disk space.

n no: means do not compile this capability at all.

3.3 Compiling

make targets

• Here are the steps to compile the base kernel image:

make dep: only needed in pre 2.6 kernels, not in 2.6 kernels

make clean: removes old object files; important if the source has been compiled previously make bzImage: builds the kernel image file. You will find it in the location arch/

    i386/boot/bzImage

• There are alternatives that I suggest you avoid, such as:

make zImage  Documentation/i386/boot.txt says this is deprecated af-

ter 2.4.0. For a very small kernel, loaded into low memory.

make zlilo attempts to install the kernel directly using lilo

make zdisk to create a bootable floppy.

• It is simplest (to me) to use make bzImage and copy the kernel file to wherever you want it.

make targets for the modules

make modules: builds the kernel modules. Takes a while on a slow machine, especially if you have enabled many kernel modules

sudo make modules_install: install the modules under /lib/modules/⟨kernel-name⟩, where ⟨kernel-name⟩ is determined by how you edited the variables at the top of the main Makefile

Other make targets

make mrproper: Clean the kernel source tree completely, to almost pristine condition. This will also delete .config. (make distclean slightly cleaner).

• Some people say the name means something highly and deeply technical (maintainer proper)

• … but Linus says it’s a cleaning fluid (German version of Mr Clean): http://www.

alphalinux.org/archives/axp-list/1996/October1996/1237.html
On Tue, 22 Oct 1996, Marc Singer wrote:
>
> > > What is mrproper? I've been wondering this for a long time.
> >
> > mrproper clears out all the config preferences.
>
> Yes, but what does it represent? Mr. Proper?
There was a silly cleaning cluid commercial over here in Finland
a few years ago with a particularly annoying jingle. "Mr Proper"
is/was the name of the cleaning fluid.
Sorry about that,
Linus

4. Installing the kernel  7

4 Installing the kernel

4.1 Installing the kernel itself

Installing into /boot

• All these files should have a name containing the version that you set in the Makefile

– In the following, replace $VERSION by the value of VERSION in the Makefile,

$PATCHLEVEL by the value of PATCHLEVEL in the Makefile, …

• Manually copy it from the file arch/i386/boot/bzImage (relative to the top

level of the source tree) to /boot

– Copy it to the name

      /boot/vmlinuz-$VERSION.$PATCHLEVEL.
      $SUBLEVEL$EXTRAVERSION
• Copy the System.map file into /boot
    – Call it /boot/System.map-$VERSION.$PATCHLEVEL.$SUBLEVEL
      $EXTRAVERSION
• Copy .config to /boot as /boot/config-$VERSION.$PATCHLEVEL.
  $SUBLEVEL$EXTRAVERSION

4.2 Make the initial RAM disk file

Make the initial RAM disk file

• If you did not compile all the modules that your kernel needs to access the hard disk right into the kernel (not as modules), then you need an initial ram disk file

• Let us represent the value of the kernel version — $VERSION.$PATCHLEVEL.

  $SUBLEVEL$EXTRAVERSION

as ⟨kernel version⟩.

• On Red Hat systems you create this with a command like this:

   $ sudo mkinitrd -v /boot/initrd-⟨kernel version⟩.img ⟨kernel version⟩
   $ mkinitrd --help
   usage: mkinitrd [--version] [-v] [-f] [--preload <module>]
           [--omit-scsi-modules] [--omit-raid-modules] [--omit-lvm-modules]
           [--with=<module>] [--image-version] [--fstab=<fstab>] [--nocompress]
           [--builtin=<module>] [--nopivot] <initrd-image> <kernel-version>
           (ex: mkinitrd /boot/initrd-2.2.5-15.img 2.2.5-15)

4.3 Having Grub start the kernel

Having Grub start the kernel

   • edit grub’s configuration file /boot/grub/menu.lst or /boot/grub/grub.
     conf

• Add a new section for your kernel, telling grub about the initrd file if you need one:

title Latest kernel (2.6.13.2)
         root (hd0,0)
         kernel /boot/vmlinuz-2.6.13.2 ro root=/dev/hda1
         initrd /boot/initrd-2.6.13.2.img

Test it

• Do not remove your old kernel from /boot/grub/menu.lst before you have tested your new kernel

• Boot the new kernel on a test system and give it a good try out before you install it on your production systems

License covering this document Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005 Nick Urbanik <nicku@nicku.org> You can redistribute modified or unmodified copies of this document provided that this

copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation — either version 2 of the License or (at your option) any later version.